OpenStudy (anonymous):

The down payment for the lot you buy is $25,000. If I have only $9500, what percent of the down payment do you still need?

OpenStudy (bahrom7893):

25000-100% 9500 - x% x = 9500*100/25000 =950000/25000=950/25=38%

OpenStudy (bahrom7893):

He needs 100-38=62%

OpenStudy (anonymous):

(25-9.5)/25 and move the decimal point 2 places right.